{2a+3b=12
{5a-b=13 please help me ,,,,thank you

Answer by pwac(253)

2a+3b=12-----(1)
5a-b=13------(2)re-arrangeto get
-b=13+5a
b=-13-5a substitute b in(1)
2a+3(-13-5a)=12 multiply bracket by 3
2a-39-15a=12 sum a-terms and add 39 to both sides
17a=51 divide both sides by 17
a=3 substitute this value of a into (1)
2(3)+3b=12
6+3b=12 subtract 6 from both sides
3b=6 divide both sides by3
b=2
a=3,b=2 substitute these into either equation to check
Pete